# Evaluating the Adversarial Robustness of Retrieval-Based In-Context Learning for Large Language Models

### v2

@@ Abstract
With the emergence of large language models, such as LLaMA and OpenAI GPT-3, In-Context Learning (ICL) gained significant attention due to its effectiveness and efficiency. However, ICL is very sensitive to the choice, order, and verbaliser used to encode the demonstrations in the prompt. Retrieval-Augmented ICL methods try to address this problem by leveraging retrievers to extract semantically related examples as demonstrations. While this approach yields more accurate results, its robustness against various types of adversarial attacks, including perturbations on test samples, demonstrations, and retrieved data, remains under-explored. Our study reveals that retrieval-augmented models can enhance robustness against test sample attacks, outperforming vanilla ICL with a 4.87% reduction in Attack Success Rate (ASR); however, they exhibit overconfidence in the demonstrations, leading to a 2% increase in ASR for demonstration attacks. Adversarial training can help improve the robustness of ICL methods to adversarial attacks; however, such a training scheme can be too costly in the context of LLMs. As an alternative, we introduce an effective training-free adversarial defence method, DARD, which enriches the example pool with those attacked samples. We show that DARD yields improvements in performance and robustness, achieving a 15% reduction in ASR over the baselines. Code and data are released to encourage further research [^1] https://github.com/simonucl/adv-retreival-icl.
@@ 1 Introduction
Large Language models (LLMs) are revolutionising the field of NLP (Brown et al., 2020; Wei et al., 2022; Touvron et al., 2023). However, this also raises concerns about their robustness and trustworthiness. Although significant efforts are put into aligning LLMs for safety purposes, recent works still show LLMs can be vulnerable to adversarial inputs and jailbreaking (Zou et al., 2023; Xu et al., 2023b; Zeng et al., 2024). As a result, it is essential to understand LLM security properties to guide the direction of LLMs that are secure and robust. In this paper, we focus on examining a common LLM inference method, namely In-Context Learning (ICL) and its variants, on how they might be susceptible to perturbations and their robustness against adversarial attacks.
@@
We address (in Section 4) the following question: How sensitive are vanilla and retrieval-based ICL methods to perturbations to their test samples or demonstration instances, which is particularly important given that the primary motivation of ICL is to facilitate few-shot learning without training; should any methods exhibit excessive sensitivity, this would limit their applicability. To investigate this problem, we employ commonly used perturbation attacks—namely character-level, word-level, and sentence-level attacks. Furthermore, we propose new attack methods that target the demonstration instances and datastore examples. Our findings indicate that retrieval-based ICL methods show greater robustness against test-sample attacks; however, they can be more vulnerable when their demonstrations are adversarially perturbed. Detailed ablation studies further reveal that this vulnerability to adversarial attacks persists in newer models, such as Mistral (Jiang et al., 2023) and Gemma (Gemma Team et al., 2024). Many of these attacks can be transferable, both from the same family (LLaMA-2-7B $\rightarrow$ LLaMA-2-13B, 70B) or different family (LLaMA-2-7B $\rightarrow$ Mistral-7B). We showed Mixture of Expert (MoE) models still posed a similar adversarial threat to its dense version. This encourages further investigation on enhancing model robustness.
We further explore the advantages of utilising retrieval-based models to improve robustness (in Section 5), considering that existing defence strategies often rely on an adversarial training process (Li et al., 2021; Si et al., 2021). However, this process is memory intensive in the context of LLMs. Instead, we show that retrieval-based ICL methods, combined with adversarially augmented examples, yield more robust results than no augmentation methods while maintaining their predictive accuracy on in-distribution evaluation sets. Therefore, we propose DARD, a training-free adversarial defence method which leverages adversarially perturbed samples to augment the retrieval pools, potentially improving the models’ robustness. The paper is summarized in the Figure 1.
@@ 2 Related Work
Several works studied this problem and the main finding is that LMs struggle to answer correctly and factually under attacks and propose their defence via label smoothing or prompt tuning (Yang et al., 2022; Raman et al., 2023). It is important to note, however, that these experiments were mainly conducted using relatively small LMs such as BERT, RoBERTa. Recently, several works demonstrated LLM can still be vulnerable to these type of attacks, by showing the vulnerability of attacking the task instructions (PromptBench (Zhu et al., 2023)), or perform adversarial attacks on zero-shot settings and investigate the transferability of attacks from open-sourced models (i.e. Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023)) to close-sourced (i.e. GPT-3.5 and GPT-4) (DecodingTrust (Wang et al., 2023a)). However, previous study (Gu et al., 2022) has shown that the LLMs are less sensitive to prompt / instruction variation when few-shot examples are provided in context (a.k.a ICL). However, there is still a lack of systematic analysis on this case which is the main contribution of this paper. Wang et al. (2023b) shows that attacking demonstrations can still be an effective way for adversarial attacks where LLMs are still vulnerable. Therefore, we aim to extend and evaluate the adversarial attack and defence performance on retrieval-based models.
@@
The existing literature on ICL analysis can be broadly divided into two streams, each focusing on different aspects. The first stream explores the influencing factors of ICL based on input perturbation, such as the order (Min et al., 2022), the formatting (Yoo et al., 2022; Lu et al., 2022) and the selection of the demonstration (Liu et al., 2022). Designing proper demonstration construction strategies (Ye et al., 2023; Li et al., 2023) could bring clear boosts to the ICL performance, while Chen et al. (2023) analyses the sensitivity of ICL via perturb on instructions and example ordering, and our work is a complement of their work on sensitivity.
This paper provides a different view of when different ICL methods fail and when retrieval-based ICL can be a better approach for robustness.
@@
Retrieval-based LLM has shown great success in lifting LLM performance and faithfulness and reducing hallucination, where the idea is to gather related articles or demonstrations from a collection of data. Specifically, we are interested in whether they can be used to improve the robustness against perturbations (Lin & Lee, 2024). Meanwhile, many papers also focus on analysing the robustness of RAG methods on irrelevant contexts (Shi et al., 2023; Asai et al., 2024), in which we adopt similar methods in our context for analysing how retrieval-based ICL being sensitive to irrelevant contexts.
@@ 3.2 Attack Methods
One of our objectives is to analyse the sensitivity of retrieval-based ICL methods to adversarial perturbations. Therefore, we consider Irrelevant Context Attacks (Irr.), which work by contaminating the demonstration pools with irrelevant contexts. Our Irrelevant Context Attacks extended Min et al. (2022) — we replace 50% of the original demonstrations in the training data with Out-of-Distrubtion (OoD) examples. Further details and analyses are available in Appendix A.3.
@@ 4.2 Results
4.3 Ablation Study
Figure 3: Attack Success Rate (%) for adversarial attacks across various models, based on experiments conducted on the RTE dataset with 8-shot demonstrations. The results are based on R-ICL and involve the aggregation of results from retrievers: BM25, SBERT, and Instructor.
@@ A.1 Details about the ICL methods
A.2 Templates and verbaliser for datasets
@@ A.3 Attack Details
    TextFooler: We also evaluated another popular word-level method, TextFooler (Jin et al., 2020), which is similar principle to TextBugger. However, TextFooler greedily searches for a large number of nearby words in the embedding space for each word, as long as they meet some constraints on embedding similarity and sentence quality. Additional constraints require the replacement words to match the Part-Of-Speech (POS) of the original words.
@@
Examples of each attack are demonstrated in Table 5-11. Code will be released to facilitate further research in this field.
A.4 Reproducability
@@ B.1 Main results
We include the complete results breakdown in Table 12-17. The results shown in those tables are the Benign accuracy and Attack accuracy.

### v3

# Evaluating and Safeguarding the Adversarial Robustness of Retrieval-Based In-Context Learning

@@ Abstract
With the emergence of large language models, such as LLaMA and GPT, In-Context Learning (ICL) gained significant attention due to its effectiveness and efficiency. However, ICL is very sensitive to the choice, order, and verbaliser used to encode the demonstrations in the prompt. Retrieval-Augmented ICL methods try to address this problem by leveraging retrievers to extract semantically related examples as demonstrations. While this approach yields more accurate results, its robustness against various types of adversarial attacks, including perturbations on test samples, demonstrations, and retrieved data, remains under-explored. Our study reveals that retrieval-augmented models can enhance robustness against test sample attacks, outperforming vanilla ICL with a 4.87% reduction in Attack Success Rate (ASR); however, they exhibit overconfidence in the demonstrations, leading to a 2% increase in ASR for demonstration attacks. Adversarial training can help improve the robustness of ICL methods to adversarial attacks; however, such a training scheme can be too costly in the context of LLMs. As an alternative, we introduce an effective training-free adversarial defence method, DARD, which enriches the example pool with those attacked samples. We show that DARD yields improvements in performance and robustness, achieving a 15% reduction in ASR over the baselines. Code and data are released to encourage further research [^1] https://github.com/simonucl/adv-retreival-icl.
@@ 1 Introduction
Large Language models (LLMs) are revolutionising the field of NLP (Brown et al., 2020; Wei et al., 2022; Touvron et al., 2023). However, this also raises concerns about their robustness (Xu et al., 2023c). Although significant efforts are put into aligning LLMs for safety purposes, recent works still show LLMs can be vulnerable to adversarial inputs and jailbreaking (Zou et al., 2023; Xu et al., 2023b; Zeng et al., 2024; Longpre et al., 2024). As a result, it is essential to understand LLM security properties to guide the direction of LLMs that are secure and robust. In this paper, we focus on examining a common LLM inference method, namely In-Context Learning (ICL) and its variants, on how they might be susceptible to perturbations and their robustness against adversarial attacks.
@@
We address (cf. Section 4) the following question: How sensitive are vanilla and retrieval-based ICL methods to perturbations to their test samples or demonstration instances, which is particularly important given that the primary motivation of ICL is to facilitate few-shot learning without training; should any methods exhibit excessive sensitivity, this would limit their applicability. To investigate this problem, we employ commonly used perturbation attacks—namely character-level, word-level, and sentence-level attacks. Furthermore, we propose new attack methods that target the demonstration instances and datastore examples. Our findings indicate that retrieval-based ICL methods show greater robustness against test-sample attacks; however, they can be more vulnerable when their demonstrations are adversarially perturbed. Detailed ablation studies further reveal that this vulnerability to adversarial attacks persists in newer models, such as Mistral (Jiang et al., 2023) and Gemma (Gemma Team et al., 2024). Many of these attacks can be transferable, both from the same family (LLaMA-2-7B $\rightarrow$ LLaMA-2-13B, 70B) or different family (LLaMA-2-7B $\rightarrow$ Mistral-7B). We showed Mixture of Expert (MoE) models still posed a similar adversarial threat to its dense version. This encourages further investigation on enhancing model robustness.
We further explore the advantages of utilising retrieval-based models to improve robustness (cf. Section 5), considering that existing defence strategies often rely on an adversarial training process (Li et al., 2021; Si et al., 2021). However, this process is memory intensive in the context of LLMs. Instead, we show that retrieval-based ICL methods, combined with adversarially augmented examples, yield more robust results than no augmentation methods while maintaining their predictive accuracy on in-distribution evaluation sets. Therefore, we propose DARD, a training-free adversarial defence method which leverages adversarially perturbed samples to augment the retrieval pools, potentially improving the models’ robustness. Figure 1 presents an overview of the paper.
@@ 2 Related Work
Several works studied this problem and the main finding is that LMs struggle to answer correctly and factually under attacks and propose their defence via label smoothing or prompt tuning (Yang et al., 2022; Raman et al., 2023). It is important to note, however, that these experiments were mainly conducted using relatively small LMs such as BERT, RoBERTa. Recently, several works demonstrated LLM can still be vulnerable to these type of attacks, by showing the vulnerability of attacking the task instructions (PromptBench (Zhu et al., 2023)), or performing adversarial attacks on zero-shot settings and investigating the transferability of attacks from open-sourced models (i.e. Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023)) to close-sourced (i.e. GPT-3.5 and GPT-4) (DecodingTrust (Wang et al., 2023a)). However, previous study (Gu et al., 2022) has shown that the LLMs are less sensitive to prompt / instruction variation when few-shot examples are provided in context (a.k.a ICL). However, there is still a lack of systematic analysis on this case which is the main contribution of this paper. Wang et al. (2023b) shows that attacking demonstrations can still be an effective way for adversarial attacks where LLMs are still vulnerable. Therefore, we aim to extend and evaluate the adversarial attack and defence performance on retrieval-based models.
@@
The existing literature on ICL analysis can be broadly divided into two streams, each focusing on different aspects. The first stream explores the influencing factors of ICL based on input perturbation, such as the order (Min et al., 2022), the formatting (Yoo et al., 2022; Lu et al., 2022) and the selection of the demonstration (Liu et al., 2022). Designing proper demonstration construction strategies (Ye et al., 2023; Li et al., 2023) could bring clear boosts to the ICL performance, while Chen et al. (2023) analyses the sensitivity of ICL via perturb on instructions and example ordering, and our work is a complement of their work on sensitivity. This paper provides a different view of when different ICL methods fail and when retrieval-based ICL can be a better approach for robustness.
@@
Retrieval-based LLM has shown great success in lifting LLM performance and faithfulness and reducing hallucination (Pan et al., 2023), where the idea is to gather related articles or demonstrations from a collection of data (Huang et al., 2023). Specifically, we are interested in whether they can be used to improve the robustness against perturbations (Lin & Lee, 2024). Meanwhile, many papers also focus on analysing the robustness of RAG methods on irrelevant contexts (Shi et al., 2023; Asai et al., 2024), in which we adopt similar methods in our context for analysing how retrieval-based ICL being sensitive to irrelevant contexts.
@@ 3.2 Attack Methods
One of our objectives is to analyse the sensitivity of retrieval-based ICL methods to adversarial perturbations. Therefore, we consider Irrelevant Context Attacks (Irr.), which work by contaminating the demonstration pools with irrelevant contexts. Our Irrelevant Context Attacks extended Min et al. (2022) — we replace 50% of the original demonstrations in the training data with Out-of-Distrubtion (OoD) examples.
Further details and analyses about attack methods are available in Appendix A.4.
@@ 4.2 Results
4.3 Attack on Model Variants
Figure 3: Attack Success Rate (%) for adversarial attacks across various models, based on experiments conducted on the RTE dataset with 8-shot demonstrations. The results are based on R-ICL and the mean attack success rate among the three retrievers we used: BM25, SBERT, and Instructor.
@@ 6 Summary
7 Limitations
We view our work as an early exploration of the adversarial robustness of Retrieval-based in-context learning. Here, we only show how different types of attack can be applied to, but not why. Given the standard use of retrieval and in-context learning as inference methods, it is crucial to understand its deficiency. Future work should focus on certified robustness (Zhao et al., 2022; Xiang et al., 2024), which analyzes with theoretical proof the extent to which the model is unaffected by adversarial attacks.
Acknowledgement
Computations described in this research were supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh. We thank Giwon Hong and Yu Zhao for the insightful discussion at the beginning of this research work.
@@ A.1 Details about the ICL methods
A.2 The statistics of datasets
Datasets No. of Labels Train/Test Size Avg. Length
SST-2 2 67348/872 23.07
RTE 2 2489/277 96.03
MR 2 8530/1066 37.76
CR 2 3394/376 30.80
MNLI-mm 3 8832/982 55.90
TREC 6 5452/500 19.59
Table 3: Datasets Information
The table 3 shows the number of categories for each classification dataset, the number of examples in the training and test sets, and the average sentence length.
A.3 Templates and verbaliser for datasets
@@ A.4 Attack Details
    TextFooler: We also evaluated another popular word-level method, TextFooler (Jin et al., 2019), which is similar principle to TextBugger. However, TextFooler greedily searches for a large number of nearby words in the embedding space for each word, as long as they meet some constraints on embedding similarity and sentence quality. Additional constraints require the replacement words to match the Part-Of-Speech (POS) of the original words.
@@
Examples of each attack are demonstrated in Table 6-12. Code will be released to facilitate further research in this field.
A.5 Reproducability
@@ B.1 Main results
We include the complete results breakdown in Table 13-18. The results shown in those tables are the Benign accuracy and Attack accuracy.